# UltraEdit: Instruction-based Fine-Grained Image Editing at Scale

### v1

@@ Abstract
This paper presents UltraEdit, a large-scale (~4M editing samples), automatically generated dataset for instruction-based image editing. Our key idea is to address the drawbacks in existing image editing datasets like InstructPix2Pix [10] and MagicBrush [59], and provide a systematic approach to producing massive and high-quality image editing samples. UltraEdit offers several distinct advantages: 1) It features a broader range of editing instructions by leveraging the creativity of large language models (LLMs) alongside in-context editing examples from human raters; 2) Its data sources are based on real images, including photographs and artworks, which provide greater diversity and reduced bias compared to datasets solely generated by text-to-image models; 3) It also supports region-based editing, enhanced by high-quality, automatically produced region annotations. Our experiments show that canonical diffusion-based editing baselines trained on UltraEdit set new records on MagicBrush and Emu-Edit benchmarks. Our analysis further confirms the crucial role of real image anchors and region-based editing data. The dataset, code, and models will be made public.
@@ 2.1 Dataset Formation
where $M$ is $I_{m}$ in the size of latent space, $DM(\cdot)$ denotes the diffusion model. In a nutshell, we alternate between regular diffusion and inpainting only within the mask region to guide the generation within the given region while avoiding edge artifacts. This pipeline is compatible with P2P control and the SDXL-Turbo backbone and it takes 3-7 diffusion steps to produce a target image. Further details can be found in Appendix A.3.
@@ 3.4 Qualitative Evaluation
In Figure 4, we provide some qualitative examples of different editing tasks on Emu Edit Test and multi-turn editing on MagicBrush. More examples can be found in Appendix D. Our main observations are: 1) most baselines, especially Emu Edit, tend to overedit the image, i.e. creating content while ignoring the context of the source image. For example, Emu Edit makes crude modifications by setting the entire person’s body as blank. We attribute these shortcomings to the biases introduced during the construction of Emu Edit’s training data; 2) Many baselines also fail to generalize to novel instructions, e.g., blurring, adding special style, etc.; 3) For multi-turn editing, even the MagicBrush baselines cannot complete the long-term editing coherently, while our model, even without editing region input, can strictly follow the instruction, e.g. one person, one bench, etc, and reaches the best results with region information, which confirms the effectiveness of region-based data in UltraEdit.
@@ 3.5 Insights and Analysis
Real Image Anchors. We conduct an ablation study to verify the effectiveness of incorporating real images as anchors during data generation. We train the editing model using two distinct datasets: free-form image editing data from UltraEdit and the editing data generated using the same pipeline of UltraEdit without using real images as anchors, i.e. identical to the data creation pipeline of InstructPix2Pix [10]. The models were trained on data volumes of 450K, 1M, and 1.5M for both datasets. The results can be found in Table 6. Our key observations: 1) Dataset generated with real image anchors generally leads to better models across all three scales; 2) The scaling effect only presents when real image anchors are adopted. We hypothesize that datasets without them could suffer from more severe image biases and therefore hinder the effect of further scaling up. In Appendix D.3, we demonstrate more qualitative results comparing the image editing generation method with and without using real images as anchors.
@@
Free-from vs. Region-based Editing. We then explore the impact of incorporating region-based editing data during model training. We experiment with varying amounts of free-form editing data, ranging from 200K to 400K instances, and region-based editing data, ranging from 30K to 90K instances. We demonstrate the performances on Emu Edit Test of models trained on the possible volume combinations of these two types of data in Figure 5. It can be seen that: 1) Region-based editing data, despite its relatively smaller scale, can help with free-form editing tasks; 2) While scaling free-form editing data has a significant impact on the performance of region-based editing, we need to ensure a considerable volume of region-based data to ensure peak results. In Appendix D.4, we demonstrate more qualitative results with region input and the model exhibits significantly more precise operations for background and localized edits, validating the effectiveness of the incorporation of the region-based editing.
@@ 4 Related Work
Image Editing via Generation.  Editing real photos according to specific instructions has long been a notable task in image processing [15, 31, 60, 46, 37]. Powerful Large-scale diffusion models have significantly facilitated text-based image editing [23, 47, 27, 13, 32]. SDEdit [33] applies noise to the guidance image at an intermediate diffusion step and then denoises using the target description. Prompt-to-Prompt [21] injecting the input caption attention maps to the target ones. Null-Text Inversion [34] inverts the source image to the null-text embedding for editing, eliminating the need for original captions. GLIDE [36] and Imagen Editor [56] fine-tuning the model to take channel-wise concatenation of the input image and mask. Blended Diffusion [5] blends the input image in the unmasked regions in the diffusion step. Meanwhile, instruction-based image editing has been introduced as a user-friendly method for image editing. Instructpix2pix [10] and HIVE [61] both are trained on generated editing data to handle user-written instructions during inference. MagicBrush [59] creates a manually-annotated dataset for fine-tuning Instructpix2pix. Emu Edit [51], trained on 10 million proprietary multi-task data, demonstrates state-of-the-art performance.
Image Editing Dataset.  Table 1 compares various image editing datasets, revealing that high-quality data are scarce and challenging to obtain. The largest dataset only contains around 300,000 samples. EditBench [57] is manually curated with only 240 examples. MagicBrush [59]manually annotated by hearing online labors using DALL-E 2 [42] for crafting dataset, also limited in dataset size. For automatically annotated datasets, InstructPix2Pix utilizes prompt-to-prompt [21] to generate data pairs based on the captions form LAION-Aesthetics [50] and edited captions generated by GPT-3 [11], but biases in the generative model and insufficient information in web-crawled captions make the generated image samples fail to represent the editing instructions, as shown in the Table 1. The the limited scope of the image instructions types of InstructPix2Pix further limits the utility of the dataset. HQ-Edit[22] uses advanced models like GPT-4 [2, 1] and DALL-E 3 [6] for generating image editing pairs, but may fail to preserve fine-grained details and realism in the target image.
@@ A.1 Collection of High-Quality Image Caption Data
We gathered data from various public data sources with image caption data, as illustrated in Table 7, which includes diverse images with either manually annotated captions or detailed captions generated by advanced image captioning models [12, 49]. After filtering out images with excessively long or short captions, our collection amounted to 1.6 million high-quality image-caption pairs, which will be used for edit instruction generation and subsequent image generation.
@@ A.2 Instruction and Caption Generation
To obtain high-quality editing instructions, we introduce a pipeline that combines human raters and language models for generating edit instructions and corresponding captions for subsequent image generation. Firstly, we use language models to expand manually crafted edit examples to a set of 100,000 examples, as shown in Table 8. These examples serve as in-context learning examples to help the language models grasp the understanding of editing styles and requirements, enabling them to generate suitable editing instructions and corresponding edited caption. The query prompt is illustrated in Table 9.
@@ B.1 Comparison with other dataset
(a) Add
(b) Change Color
(c) Change Global
(d) Change Local
(e) Others
(f) Replace
(g) Transform Global
(h) Transform Local
(i) Turn
@@ C.2 Details on Benchmarks and Metrics
Emu Edit Test aims for reducing bias of the annotator defined dataset and reach higher diversity. It contains the devise relevant, creative, and challenging instructions and high quality captions that capture both important elements in the image for source and target images, without any ground truth images. Consequently, consistent with the Emu Edit [51], we utilize the L1 distance, CLIP image similarity and DINO similarity between the source images and edited images to measure the the model‘s ability of preserving elements from the source image. Also, we use the CLIP text-image similarity between edited image and output caption and the CLIP text-image direction similarity(CLIPdir) to measure the instruction following ability of the model. Specifically, the CLIPdir measures agreement between change in caption embedding and the change in image embedding. Despite the Emu Edit Test eliminating bias and overfitting at the image level by not providing ground-truth images, the evaluation metrics still implicitly measure the model’s editing ability.
@@ F.1 Motivation
This dataset was created by Haozhe Zhao (Peking University), Xiaojian Ma (BIGAI), Liang Chen (Peking University), Shuzheng Si (Peking University), Rujie Wu (Peking University), Kaikai An (Peking University), Peiyu Yu (UCLA), Minjia Zhang (UIUC), Qing Li (BIGAI), and Baobao Chang (Peking University).

### v2

@@ Abstract
This paper presents UltraEdit, a large-scale (~4M editing samples), automatically generated dataset for instruction-based image editing. Our key idea is to address the drawbacks in existing image editing datasets like InstructPix2Pix [10] and MagicBrush [71], and provide a systematic approach to producing massive and high-quality image editing samples. UltraEdit offers several distinct advantages: 1) It features a broader range of editing instructions by leveraging the creativity of large language models (LLMs) alongside in-context editing examples from human raters; 2) Its data sources are based on real images, including photographs and artworks, which provide greater diversity and reduced bias compared to datasets solely generated by text-to-image models; 3) It also supports region-based editing, enhanced by high-quality, automatically produced region annotations. Our experiments show that canonical diffusion-based editing baselines trained on UltraEdit set new records on MagicBrush and Emu-Edit benchmarks. Our analysis further confirms the crucial role of real image anchors and region-based editing data. The dataset, code, and models are available in github.com/pkunlp-icler/UltraEdit.
@@ 2.1 Dataset Formation
where $M$ is $I_{m}$ in the size of latent space, $DM(\cdot)$ denotes the diffusion model. In a nutshell, we alternate between regular diffusion and inpainting only within the mask region to guide the generation within the given region while avoiding edge artifacts, as illustrated in Figure 16. This pipeline is compatible with P2P control and the SDXL-Turbo backbone and it takes 3-7 diffusion steps to produce a target image. Further details can be found in Appendix A.3.
@@ 3.3 Main Result II: Scaling Effect of UltraEdit on Emu Edit Test
Figure 4: Qualitative comparison of images generated by our data generation pipeline with other zero-shot image editing methods.
@@ 3.4 Qualitative Evaluation
Data Generation.
In Figure 4, we present qualitative examples generated by our data generation pipeline alongside various image editing methods for comparison. We evaluate our pipeline against several baselines, including image inversion methods like Null-text Inversion [42] and PnP Inversion [27]; inpainting methods such as BrushNet [28] and PowerPaint [76]; and zero-shot image editing methods like InfEdit [69] and MasaCtrl [12]. Our data generation pipeline, leveraging SDXL-turbo, requires only 1-4 steps to produce a sample, which is significantly faster than other methods by magnitudes. Importantly, this efficiency does not compromise the quality. Our pipeline demonstrates competitive performance against other baselines, highlighting the quality of  UltraEdit.
Image Editing.
In Figure 5, we provide some qualitative examples of different editing tasks on Emu Edit Test and multi-turn editing on MagicBrush generated by the Stable Diffusion v1.5 trained on  UltraEdit. More examples can be found in Appendix D. Our main observations are: 1) most baselines, especially Emu Edit, tend to overedit the image, i.e. creating content while ignoring the context of the source image. For example, Emu Edit makes crude modifications by setting the entire person’s body as blank. We attribute these shortcomings to the biases introduced during the construction of Emu Edit’s training data; 2) Many baselines also fail to generalize to novel instructions, e.g., blurring, adding special style, etc.; 3) For multi-turn editing, even the MagicBrush baselines cannot complete the long-term editing coherently, while our model, even without editing region input, can strictly follow the instruction, e.g. one person, one bench, etc, and reaches the best results with region information, which confirms the effectiveness of region-based data in UltraEdit.
@@ 3.5 Insights and Analysis
Real Image Anchors.
We conduct an ablation study to verify the effectiveness of incorporating real images as anchors during data generation. We train the editing model using two distinct datasets: free-form image editing data from UltraEdit and the editing data generated using the same pipeline of UltraEdit without using real images as anchors, i.e. identical to the data creation pipeline of InstructPix2Pix [10]. The models were trained on data volumes of 450K, 1M, and 1.5M for both datasets. The results can be found in Table 6. Our key observations: 1) Dataset generated with real image anchors generally leads to better models across all three scales; 2) The scaling effect only presents when real image anchors are adopted. We hypothesize that datasets without them could suffer from more severe image biases and therefore hinder the effect of further scaling up. In Appendix D.3, we demonstrate more qualitative results comparing the image editing generation method with and without using real images as anchors.
@@
Free-from vs. Region-based Editing.
We then explore the impact of incorporating region-based editing data during model training. We experiment with varying amounts of free-form editing data, ranging from 200K to 400K instances, and region-based editing data, ranging from 30K to 90K instances. We demonstrate the performances on Emu Edit Test of models trained on the possible volume combinations of these two types of data in Figure 6. It can be seen that: 1) Region-based editing data, despite its relatively smaller scale, can help with free-form editing tasks; 2) While scaling free-form editing data has a significant impact on the performance of region-based editing, we need to ensure a considerable volume of region-based data to ensure peak results. In Appendix D.4, we demonstrate more qualitative results with region input and the model exhibits significantly more precise operations for background and localized edits, validating the effectiveness of the incorporation of the region-based editing.
@@ 4 Related Work
Image Editing via Generation.
Editing real photos according to specific instructions has long been a notable task in image processing [21, 16, 39, 72, 54, 45]. Powerful Large-scale diffusion models have significantly facilitated text-based image editing [29, 55, 33, 14, 40]. SDEdit [41] applies noise to the guidance image at an intermediate diffusion step and then denoises using the target description. Prompt-to-Prompt [24] injecting the input caption attention maps to the target ones. Null-Text Inversion [42] inverts the source image to the null-text embedding for editing, eliminating the need for original captions. Plug-and-Play [64] incorporates spatial features besides attention maps for better global image editing. Imagic [29] supports complex textual instructions via text embedding optimization and model fine-tuning. GLIDE [44] and Imagen Editor [67] fine-tuning the model to take channel-wise concatenation of the input image and mask. Blended Diffusion [5] blends the input image in the unmasked regions in the diffusion step. Meanwhile, instruction-based image editing has been introduced as a user-friendly method for image editing. Instructpix2pix [10] and HIVE [73] both are trained on generated editing data to handle user-written instructions during inference. MagicBrush [71] creates a manually-annotated dataset for fine-tuning Instructpix2pix. Emu Edit [60], trained on 10 million proprietary multi-task data, demonstrates state-of-the-art performance.
Image Editing Dataset.
Table 1 compares various image editing datasets, revealing that high-quality data are scarce and challenging to obtain. The largest dataset only contains around 300,000 samples. EditBench [68] is manually curated with only 240 examples. MagicBrush [71]manually annotated by hearing online labors using DALL-E 2 [50] for crafting dataset, also limited in dataset size. For automatically annotated datasets, InstructPix2Pix utilizes prompt-to-prompt [24] to generate data pairs based on the captions form LAION-Aesthetics [59] and edited captions generated by GPT-3 [11], but biases in the generative model and insufficient information in web-crawled captions make the generated image samples fail to represent the editing instructions, as shown in the Table 1. The the limited scope of the image instructions types of InstructPix2Pix further limits the utility of the dataset. HQ-Edit[26] uses advanced models like GPT-4 [2, 1] and DALL-E 3 [6] for generating image editing pairs, but may fail to preserve fine-grained details and realism in the target image.
@@ 5 Conclusion
Acknowledgement
We extend our heartfelt gratitude to the anonymous reviewers whose dedication and insightful feedback have significantly enhanced the caliber of this paper. Their constructive critiques and valuable suggestions were instrumental in refining our work. Additionally, we are deeply appreciative of the Program Chairs and Area Chairs for their meticulous handling of our submission and for their comprehensive and invaluable feedback. Their guidance has been pivotal in elevating the quality of our research.
This work is supported by the National Science Foundation of China under Grant No.61936012 and 61876004.
@@ A.1 Collection of High-Quality Image Caption Data
Like previous works [34, 13, 71, 75], we gathered source data from various public data sources with image caption data, as illustrated in Table 7, which includes diverse images with either manually annotated captions or detailed captions generated by advanced image captioning models [13, 58]. After filtering out images with excessively long or short captions, our collection amounted to 1.6 million high-quality image-caption pairs, which will be used for edit instruction generation and subsequent image generation.
@@ A.2 Instruction and Caption Generation
To obtain high-quality editing instructions, we introduce a pipeline that combines human raters and language models for generating edit instructions and corresponding captions for subsequent image generation. Large language models have demonstrated remarkable abilities in various areas, such as agents [18, 25, 61, 35] and tool uage [57]. In our practice, we utilize the LLM to generate suitable image editing instructions. Firstly, we use language models to expand manually crafted edit examples to a set of 100,000 examples, as shown in Table 8. These examples serve as in-context learning examples to help the language models grasp the understanding of editing styles and requirements, enabling them to generate suitable editing instructions and corresponding edited caption. The query prompt is illustrated in Table 9.
@@ Appendix B Statistics of UltraEdit
(a) Add
(b) Change Color
(c) Change Global
(d) Change Local
(e) Others
(f) Replace
(g) Transform Global
(h) Transform Local
(i) Turn
@@ C.2 Details on Benchmarks and Metrics
Emu Edit Test aims for reducing bias of the annotator defined dataset and reach higher diversity. It contains the devise relevant, creative, and challenging instructions and high quality captions that capture both important elements in the image for source and target images, without any ground truth images. Consequently, consistent with the Emu Edit [60], we utilize the L1 distance, CLIP image similarity and DINO similarity between the source images and edited images to measure the the model‘s ability of preserving elements from the source image. Also, we use the CLIP text-image similarity between edited image and output caption and the CLIP text-image direction similarity(CLIPdir) to measure the instruction following ability of the model. Specifically, the CLIPdir measures agreement between change in caption embedding and the change in image embedding.Since the Emu Edit [60] does not specify the versions of the CLIP and DINO models used for the metric, we adopted the settings utilized by MagicBrush to maintain alignment with other benchmarks. Specifically, the versions are ViT-B/32 for CLIP and dino_vits16 for DINO embeddings. We ensure consistency by rerunning all results of different methods on Emu Edit benchmark. Additionally, there are known issues with the quality of the benchmark, wherein some image-caption pairs appear incorrect. These issues include placeholder captions (e.g., ’a train station in city’) or instances where source and target captions are identical. To address these problems, we simply remove the incorrect cases prior to evaluation. Despite the Emu Edit Test eliminating bias and overfitting at the image level by not providing ground-truth images, the evaluation metrics still implicitly measure the model’s editing ability.
@@ D.3 Qualitative Evaluation on Real Image Anchors
D.5 Details of the region-based image editing pipeline
Figure 16: Qualitative evaluations of the region-based image editing pipeline. Generated images Without our method exhibit noticeable artifacts along the boundaries of the original and edited regions, emphasizing pronounced border effects.
In Stage III, we apply our proposed method for generating region-based images to ensure a seamless transition between inpainted areas and the rest of the image. Initially, we analyze inaccurate masks generated by the segmentation model, as shown in Figure 17. We find these inaccuracies generally fall into a few categories: incorrect identification resulting in overly large masks, masks that are too small for effective editing, fragmented masks from segmentation failures, and fine-grained segment masks that closely resemble the original object, complicating the editing process.
To address these issues, we filter out excessively large, small, or fragmented masks. Fine-grained masks are adjusted using a soft mask, either a bounding box or contour mask. Our data circulation indicates that our methods significantly reduce artifacts and abrupt boundaries between the mask region and the remaining image. Qualitative evaluations shown in Figure 16 demonstrate the effectiveness of our approach. Images generated without our method exhibit noticeable artifacts along the boundaries of the original and edited regions, highlighting the advantages of our method.
Figure 17: The four main categories of inaccuracies generated masks.
@@ F.1 Motivation
This dataset was created by Haozhe Zhao (Peking University), Xiaojian Ma (BIGAI), Liang Chen (Peking University), Shuzheng Si (Tsinghua University), Rujie Wu (Peking University), Kaikai An (Peking University), Peiyu Yu (UCLA), Minjia Zhang (UIUC), Qing Li (BIGAI), and Baobao Chang (Peking University).
@@ F.6 Uses
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
  • Did you include the license to the code and datasets? [Yes] See supplementary.
  • Did you include the license to the code and datasets? [Yes] See supplementary.
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See supplementary.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See supplementary.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See supplementary.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See supplementary.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] See supplementary.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See supplementary.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See supplementary.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See supplementary.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]